Circular No. 7081

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA MUSCAE 1998 S. J. O’Meara, Volcano, HI, reports that prediscovery images of this

nova at m_{pg} ∼ 8.0 appear on three successive photographs taken with Fuji Sensia 400 color slide film beginning at 1998 Dec. 27.60 UT. The ∼ 10min exposures were taken with a guided 35-mm camera (+ 50-mm lens). Corrigendum. On IAUC 7079, line 7, for Dec. 1.1754 read Dec. 31.1754

MXB 1730–335 D. W. Fox and W. H. G. Lewin, Massachusetts Institute of Technology,

report on behalf of a collaboration: “We have performed a timing analysis of the first second of data from each of 31 type-I x-ray bursts from the rapid burster observed from Nov. 1996 to Feb. 1998 with the Rossi X-ray Timing Explorer (energy range 5.5–16 keV only). We find a statistically significant pair of peaks in the summed power-density spectrum, with peak centroid frequencies of 154.9 ± 0.1 and 306.6 ± 0.1 Hz, respectively. We suggest that this finding indicates a neutron-star spin period of 6.5 ms for the rapid burster.”

RX J0052.1–7319 R. C. Lamb and T. A. Prince, California Institute of Technology; D. J.

Macomb, Goddard Space Flight Center, NASA; and M. H. Finger, Marshall Space Flight Center, NASA, report their discovery of 15.3-s pulsations from the variable x-ray source RX J0052.1–7319 (1E 0050.3–7335) from analysis of 1996 data: “The pulsations are observed in both ROSAT HRI observations (0.1–2 keV) that span the interval 1996 Nov. 10.71–Dec. 9.16 UT and contemporaneous BATSE observations (20–50 keV). During this outburst, the neutron star assumed to be responsible for the pulsations is observed to be spinning up on a timescale of 38 yr with a frequency time derivative of 5.4 × 10^{−11} Hz/s. Significant power is seen in the fundamental, as well as in the second and third harmonics. The pulsed fraction in the HRI is 0.27(2). The counting rate, 0.7 counts/s, corresponds to a preliminary flux value of 2.6 × 10^{−11} erg s^{−1} cm^{−2} and a luminosity (0.1–2 keV) at an assumed distance of 60 kpc (appropriate for the SMC) of 1.1 × 10^{37} erg/s. The position of the source (Kahabka and Pietsch 1996, A.Ap. 312, 919) is α = 0^{h}52^{m}11^{s}.3, δ = −73^{o}19′13′′ (equinox 2000.0). No optical counterpart for the source is yet reported.”

1999 January 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green